A hat and a jacket together cost P1000. The jacket costs P900 more than the hat. What is the cost of the hat and the cost of the jacket?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39799)

simple arithmetic

Answer by ikleyn(53763)

.

        Solve it THIS WAY

 
   hat + jacket  = 1000

    h  +  j      = 1000

    h  + (h+900) = 1000

    2h   +  900  = 1000

        2h       = 1000 - 900

        2h       =  100

         h       = 100/2 = 50.


ANSWER.  The hat costs P 50;  the jacket costs  P 950.


        It is one way.   There is ANOTHER WAY,  if you are just familiar
        with systems of two equations in two unknowns.


From the condition, you write this system of two equations


    j + h = 1000     (1)

    j - h =  900     (2)


Add the equations. You will get then


    2j    = 1900

     j    = 1900/2 = 950.


You will get the same answer.


So,  now you know  TWO  WAYS  to solve the problem.

------------------

If you want to see a million similar  (or different)  solved problems,  look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.